OpenStudy (anonymous):

Solve for x. 4(x – b) = x A. b = 4/3x B. b = 3/4x C. x = 4/3b D. x = 3/4b

OpenStudy (anonymous):

well sense you are solving for x can you eliminate two of your choices

OpenStudy (anonymous):

Yes A and B

OpenStudy (anonymous):

x=4b/3

OpenStudy (anonymous):

Do you mean x = 4/3 b ? Joselito

OpenStudy (anonymous):

yeah but wait im seeing if its right

OpenStudy (anonymous):

Do you want to learn how to do it or just the answer

OpenStudy (anonymous):

Just the answer please

OpenStudy (anonymous):

D is the answer to me

OpenStudy (anonymous):

Distribute 4. solve. 4x - 4b = x 3x = 4b /3 x = 4/3x

OpenStudy (anonymous):

x = 4/3b

OpenStudy (anonymous):

Ok thanks